OpenStudy (anonymous):

How many photons are produced in a laser pulse of 0.628 J at 565 nm?

OpenStudy (anonymous):

wavelength = wavespeed/frequency freq = c/wavelength =3E8/550E-9 = 5.45E14 Hz photons have energy E = hf = 6.626E-34 Js * 5.45E14 Hz = 3.61 E-19 J so if it contains 4.40 mJ of energy, the pulse has: 4.40E-3/3.61E-19 = 1.22E16 photons